# Patterns vs. Patients: Evaluating LLMs against Mental Health Professionals on Personality Disorder Diagnosis through First-Person Narratives

### v1

@@ 3.2.1 Semantic Embeddings
Figure 3: MDS and UMAP projections of the semantic embeddings of diagnosis justifications.
The MDS and UMAP projections of the semantic embeddings of diagnosis justifications are visualized in Figure 3. Overall, for this analysis, MDS proved to be a more informative dimensionality-reduction method than UMAP, contributing to most of the observations summarized as follows.
@@
Llama 3.3 70B exhibits atypical semantics compared to every other agent. The distinctiveness of Llama’s semantics does not stem from undertaking vastly different diagnostic pathways than the rest of the agents, but from its comparably poor command of the Polish language. Llama’s explanations tended to contain incorrect grammar, nonexistent words, or seemingly random ,,artifact” tokens unrelated grammatically and semantically to the rest of the justifications’ linguistic content. Interestingly, this did not seem to have significantly impeded Llama’s diagnosis capabilities, as it was the ex-aequo third-best performing model (see Figure 1).
@@ 4 Discussion
Regarding the conceptualization of mental disorders, our results highlight a tension between evolving psychiatric frameworks and the persistence of traditional diagnostic classifications. While human experts performed better on the categorical than the dimensional task, models achieved relatively balanced scores across both, with a performance preference for categorical diagnosis. This suggests that the current shift toward a dimensional understanding of personality pathology APA2022; WHO2022 may be overshadowed by the legacy of earlier categorical systems, whose labels remain prevalent in everyday discourse and online corpora.
@@
The prominence of such labels in training corpora may also explain the severe underdiagnosing of NPD by models. This could stem from the negative semantic load and social stigma associated with the term ,,narcissist”, combined with the current LLM preference alignment training regimen (i.e., RLHF techniques) rewarding agreeable and non-confrontational behavior, sometimes bordering on (social) sycophancy Cheng2025; sharma2025understandingsycophancylanguagemodels. Assigning a stigmatized label to a first-person narrator—essentially ,,calling the user a narcissist”—conflicts with these preferences. Importantly, this avoidance does not reflect inability, as models showed higher dimensional recall for NPD cases, indicating sensitivity to severity without using value-laden labels. A similar aversion to stigmatizing terminology may explain the depathologizing bias observed in the GPT models family, where avoidance of psychiatric labels can downplay genuine symptoms and delay necessary treatment Semigran2015. Conversely, BPD is increasingly treated in the literature as a clinical condition rather than a pejorative label. The comparatively extensive body of BPD research relative to other disorders Blashfield2000 likely contributes to the models’ inclination to overdiagnose it.
@@
Notably, a critical divergence between humans and LLMs was observed in the assessment of confidence. Models never utilized the ,,guessing” or low-confidence options, whereas human experts frequently expressed uncertainty. This might pose a potential safety risk for clinical application. As argued by Ulmer et al. Ulmer2025, the ability of an AI to express doubt in a ,,human-like” manner is a necessary safety feature to ensure users can trust the uncertainty being conveyed. Paradoxically, although both humans and models demonstrated higher objective performance on the categorical task, they consistently reported higher subjective certainty regarding their dimensional ratings. Future studies could investigate the underlying mechanisms of this metacognitive dissociation, examining why both human and non-human agents perceive their answers regarding dimensional severity as more certain, despite their objective performance superiority in categorical classification.
@@
Furthermore, the semantic analysis reveals that although models may achieve higher accuracy, their ,,reasoning” remains fundamentally distinct from clinical practice. Human justifications were concise, cautious, and patient-centered. Model justifications, conversely, were elaborate, formulaic, and pattern-focused. Additionally, models that showed reduced performance also had more atypical semantics. This suggests a link between diagnostic accuracy and justification semantics; lower-performing models may have failed to attend to the most clinically informative aspects of the patient narratives. Llama 3.3 70B was a notable exception, achieving high diagnostic accuracy despite generating justifications riddled with linguistic artifacts, nonexistent words, and poor Polish grammar. This can be understood in terms of a multilingual generative-discriminative performance gap, in which Llama 3.3 70B demonstrates sufficient passive understanding of Polish to assign accurate diagnoses, while exhibiting markedly weaker Polish fluency when generating justifications for those decisions. Follow-up research could investigate how the performance of this model would change if its diagnostic reasoning were generated before assigning diagnostic scores, rather than as a post-hoc justification.
@@ 5 Conclusion
In this study, we provided a first-of-its-kind evaluation of state-of-the-art LLMs on raw, first-person patient data, assessed under a novel expert-informed evaluation protocol and directly compared with human specialists. Our findings indicate that current LLMs possess a surprisingly high level of diagnostic competence, largely surpassing mental health professionals in the examined context. At the same time, the models demonstrated susceptibility to bias and overconfidence. Taken together, these findings underscore the need for a collaborative human–AI framework, in which the clinical judgment and ethical oversight of experts balance the analytical consistency of models. Such integration has the potential to mitigate the respective limitations of each agent—model bias and human variability—thereby achieving higher diagnostic validity than either could attain alone.
@@
This study has several limitations. First, humans and LLMs were compared on the textual diagnostic modality only, whereas actual psychiatric assessment is inherently multimodal; clinicians rely on visual cues, affect, prosody, and interactional dynamics that are absent in text transcripts. Accordingly, the human experts in our experiment reported difficulties in relying solely on text without seeing the patient. However, given the rapid post-COVID rise of remote medicine, as well as the more recent surge in AI-driven therapeutic chatbots, the textual modality may gain prominence in the future. The second limitation is that the sample comprised only seven narratives. Although small, this sample size reflects a deliberate trade-off necessitated by the depth of the data and the substantial cognitive and time demands of human experts. The fact that this study was conducted within a specific cultural context may also be considered a limitation. This is because the diagnostic approaches of human experts and the narrative styles of patients may not be generalizable to other populations. Future research should explore whether these findings remain consistent across different cultural settings.
@@
All procedures performed in this study involving human participants were in accordance with the ethical standards of the institutional and/or national research committee and with the 1964 Helsinki Declaration and its later amendments or comparable ethical standards. Human participants signed an informed consent form by hand. The qualitative interview protocol was approved by the hospital and the Bioethics Committee of Poznan University of Medical Sciences (decision no. KB-367/23). All personal information was anonymized, ensuring full privacy of the patients’ identities.
@@ Appendix A Clinical Assessment Instruments
Levels of personality functioning and maladaptive trait domains were assessed using the Level of Personality Functioning—Brief Scale 2.0 (LPFS-BF 2.0) Lakuta2023 and the Personality Inventory for ICD-11 Cieciuch2022, ensuring comparability across individuals. LPFS-BF 2.0 scores were interpreted relative to established cut-offs Weekers2023, yielding a dimensional classification from subclinical to extreme impairment (0–48), independent of the categorical BPD/NPD labels.
Appendix B Model Prompts
@@ Appendix F Dimensionality Reduction Hyperparameters
UMAP n_components 2
n_neighbors 4
min_dist 0.1
random_state 42
Table F.2: Dimensionality reduction algorithms hyperparameter values.
Hyperparameters not present in the table retained the default values set in the algorithm implementations. The n_neighbors parameter for UMAP, which controls the balance between local and global structure, was set to a low value to accommodate the small dataset size of $N=17/16$ (with and without the human embedding). This setting prioritizes the detection of latent local substructures and family-specific clusters within the models.

### v2

@@ 3.2.1 Semantic Embeddings
Figure 3: MDS projections of the semantic embeddings of diagnosis justifications.
The MDS projections of the semantic embeddings of diagnosis justifications are visualized in Figure 3.
@@
Llama 3.3 70B exhibits atypical semantics compared to every other agent. The distinctiveness of Llama’s semantics does not stem from undertaking vastly different diagnostic pathways than the rest of the agents, but from its comparably poor command of the Polish language. Llama’s explanations tended to contain incorrect grammar, nonexistent words, or seemingly random ,,artifact” tokens unrelated grammatically and semantically to the rest of the justifications’ linguistic content. Interestingly, this did not impede Llama’s diagnosis capabilities, as it was, in fact, the ex-aequo third-best performing model (see Figure 1).
@@ 4 Discussion
Regarding the conceptualization of mental disorders, our results highlight a tension between evolving psychiatric frameworks and the persistence of traditional diagnostic classifications. While human experts performed better on the categorical than the dimensional task, models achieved relatively balanced scores across both, with a performance preference for categorical diagnosis. This suggests that the current shift toward a dimensional understanding of personality pathology American Psychiatric Association (2022); World Health Organization (2022) may be overshadowed by the legacy of earlier categorical systems, whose labels remain prevalent in everyday discourse and online corpora.
@@
The prominence of such labels in training corpora may also explain the severe underdiagnosing of NPD by models. This could potentially stem from the negative semantic load and social stigma associated with the term ,,narcissist”, combined with the current LLM preference alignment training regimen (i.e., RLHF techniques) rewarding agreeable and non-confrontational behavior, sometimes bordering on (social) sycophancy Cheng et al. (2025); Sharma et al. (2025). We hypothesize that assigning a stigmatized label to a first-person narrator—essentially ,,calling the user a narcissist”—conflicts with these preferences. Importantly, this avoidance does not reflect inability, as models showed higher dimensional recall for NPD cases, indicating sensitivity to severity without using value-laden labels. A similar aversion to stigmatizing terminology may explain the depathologizing bias observed in the GPT models family, where avoidance of psychiatric labels can downplay genuine symptoms and delay necessary treatment Semigran et al. (2015). Conversely, BPD is increasingly treated in the literature as a clinical condition rather than a pejorative label. The comparatively extensive body of BPD research relative to other disorders Blashfield and Intoccia (2000) likely contributes to the models’ inclination to overdiagnose it.
@@
Notably, a critical divergence between humans and LLMs was observed in the assessment of confidence. Models never utilized the ,,guessing” or low-confidence options, whereas human experts frequently expressed uncertainty. This might pose a potential safety risk for clinical application. As argued by Ulmer et al. Ulmer et al. (2025), the ability of an AI to express doubt in a ,,human-like” manner is a necessary safety feature to ensure users can trust the uncertainty being conveyed. Paradoxically, although both humans and models demonstrated higher objective performance on the categorical task, they consistently reported higher subjective certainty regarding their dimensional ratings. Future studies could investigate the underlying mechanisms of this metacognitive dissociation, examining why both human and non-human agents perceive their answers regarding dimensional severity as more certain, despite their objective performance superiority in categorical classification.
@@
Furthermore, the semantic analysis reveals that although models may achieve higher accuracy, their ,,reasoning” remains fundamentally distinct from clinical practice. Human justifications were concise, cautious, and patient-centered. Model justifications, conversely, were elaborate, formulaic, and pattern-focused. Additionally, models that showed reduced performance also had more atypical semantics. This suggests a link between diagnostic accuracy and justification semantics; lower-performing models may have failed to attend to the most clinically informative aspects of the patient narratives. Llama 3.3 70B was a notable exception, achieving high diagnostic accuracy despite generating justifications riddled with linguistic artifacts, nonexistent words, and poor Polish grammar. This can be understood in terms of a multilingual generative-discriminative performance gap: Llama 3.3 70B demonstrates sufficient understanding of Polish to assign accurate diagnoses, but exhibits markedly weaker Polish generation fluency when creating justifications for those decisions. Follow-up research could investigate how the model’s performance would change if its diagnostic reasoning were generated before assigning diagnostic scores, rather than as a post-hoc justification.
@@ 5 Conclusion
In this depth-first psychiatric case study, we evaluated state-of-the-art LLMs on raw, first-person patient data, assessed under a novel expert-informed evaluation protocol, and directly compared their performance with human specialists. Our observations suggest that current LLMs demonstrate a surprisingly high level of diagnostic competence, achieving higher diagnostic scores than mental health professionals in the examined context. At the same time, the models demonstrated susceptibility to bias and overconfidence. Taken together, these findings underscore the value of a collaborative human–AI framework, in which the clinical judgment and ethical oversight of experts balance the analytical consistency of models. Such integration has the potential to mitigate the respective limitations of each agent—model bias and human variability—thereby achieving higher diagnostic validity than either could attain alone.
@@
This study has several limitations. First, humans and LLMs were compared on the textual diagnostic modality only, whereas actual psychiatric assessment is inherently multimodal, which could explain the human expert’s comparatively lower performance and low inter-rater reliability in this strictly textual format. However, given the rapid post-COVID rise of remote medicine, as well as the more recent surge in AI-driven therapeutic chatbots, the textual modality may gain prominence in the future.
Second, although a sample size of seven narratives is notably small by the standards of conventional machine learning evaluations, within the context of this study, it reflects a deliberate tradeoff. The limiting factor was the immense cognitive demand of the expert evaluation task. Unlike standard clinical vignettes, the cases used in this study consisted of raw, verbatim transcripts of semi-structured autobiographical interviews drawn from a corpus exceeding 200,000 words, often resulting in over 30 pages of chaotic, non-linear narration per patient. Properly analyzing just one of these narratives required approximately five hours of expert time, as it involved careful reading, cross-referencing temporal inconsistencies, and synthesizing a coherent clinical picture from a ,,messy” life story. Consequently, securing more than 35 hours of active work from practicing, high-level experts proved logistically unfeasible due to time constraints and the significant cognitive toll. Expanding the number of narratives would have introduced a severe risk of expert fatigue; the heavy cognitive load required to review such complex qualitative data at scale would have inevitably compromised the depth and quality of the human diagnostic judgments, thereby undermining the depth-first nature of this case study.
Additionally, since diagnostic approaches of human experts are culture-specific, they may not generalize to other populations.
@@
All procedures performed in this study involving human participants were in accordance with the ethical standards of the institutional and/or national research committee and with the 1964 Helsinki Declaration and its later amendments or comparable ethical standards. Human participants signed an informed consent form by hand. The qualitative interview protocol was approved by the hospital and the Bioethics Committee of [Anonymized] University of Medical Sciences (decision no. [Anonymized]). All personal information was anonymized, ensuring full privacy of the patients’ identities.
@@ Appendix A Clinical Assessment Instruments
Levels of personality functioning and maladaptive trait domains were assessed using the Level of Personality Functioning—Brief Scale 2.0 (LPFS-BF 2.0) Łakuta et al. (2023) and the Personality Inventory for ICD-11 Cieciuch et al. (2022), ensuring comparability across individuals. LPFS-BF 2.0 scores were interpreted relative to established cut-offs Weekers et al. (2023), yielding a dimensional classification from subclinical to extreme impairment (0–48), independent of the categorical BPD/NPD labels.
Appendix B Prompts
@@ Appendix F Dimensionality Reduction Hyperparameters
Table F.2: Dimensionality reduction algorithm hyperparameter values.
Hyperparameters not present in the table retained the default values set in the algorithm implementation.